OpenStudy (zenmo):

A semicircular plate with radius 9 m is submerged vertically in water so that the top is 3 m above the surface. Express the hydrostatic force against one side of the plate as an integral and evaluate it. (Round your answer to the nearest whole number. Use 9.8 m/s2 for the acceleration due to gravity. Recall that the weight density of water is 1000 kg/m3.)

OpenStudy (zenmo):

Expressing the integral as \[2pg \int\limits_{3}^{9}[(x-3)\sqrt{81-x^2}]dx\] How would I solve this?

Parth (parthkohli):

Substitute \(x = 9 \cos \theta \).

OpenStudy (mrnood):

 Which orientation!!!